A shaft is timed to rotate 87 rev in 0.50 min. What is its angular speed in radians per second?


1
Expert's answer
2021-05-24T09:00:35-0400
"\omega=\dfrac{87\ rev}{0.5\ min}=174\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ s}\cdot\dfrac{2\pi\ rad}{1\ rev}=18.2\ \dfrac{rad}{s}."